Your firm has the following income statement items: sales of $50,250,000; income tax of $1744000; operating expenses of $8750000; cost of goods sold of $35025000; depreciation and amortization of $1365000; and interest expense of $750000. What is the amount of the firm's after tax cash flow?
Answer #1
Solution:
Firm's after tax cash flow 4,731,000
Working Notes:
Computation of after tax cash flow
A Sales 50,250,000
B Less: Cost of goods sold 35025000
C Less: operating expenses 8,750,000
D Less: Depreciation and amortization $1,365,000
E=A-B-C-D Profit Before Tax $5,110,000
F Less: income tax 1744000
G=E-F Net Income $3,366,000
H Add: Depreciation and amortization 1,365,000
I=G + H Operating Cash Flow 4,731,000
(after tax cash flow)
after tax cash flow (operating cash flow)
= Net income + Depreciation
Please feel free to ask if anything about above solution in comment section of the question.